OpenStudy (anonymous):

Mary wants to hang a mirror in her room. The mirror and frame must have an area of 7 square feet. The mirror is 2 feet wide and 3 feet long. Which quadratic equation can be used to determine the thickness of the frame, x?

OpenStudy (anonymous):

If x is the thickness of the frame, then 2x + 2 is the width and 2x + 3 is the length of the mirror with frame. Area of the frame now is 2x + 2 times 2x + 3, which should equal 7 So the equation becomes (2x + 2)(2x + 3) = 7 Simplifying it you get 4x^2 + 10x - 1 = 0

OpenStudy (anonymous):

Thanks so much, want to help with another?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

The graph of f(x) = x2 is shifted 3 units down to obtain the graph of g(x). Which of the following equations best describes g(x)?

OpenStudy (anonymous):

@Cpt.Sparz

OpenStudy (anonymous):

Is this multiple choice? Or do I have to go step by step? o:

OpenStudy (anonymous):

g(x) = (x + 3)^2 g(x) = x^2 - 3 g(x) = x^2 + 3 g(x) = (x - 3)^2

OpenStudy (anonymous):

I believe the answer is g(x) = (x - 3)^2